OpenStudy (anonymous):

what is the slope of the line containing the points (1, 3) and (-4, 0).

OpenStudy (lgbasallote):

\[\huge slope = \frac{y_2 - y_1}{x_2 - x_1}\] \[\huge \implies slope = \frac{0 - 3}{-4-1}\] got it/

OpenStudy (anonymous):

its -3/-5?

Parth (parthkohli):

Slope is supposed to be change in y over change in x, but there is a crappy formula too. \(\Rightarrow \Large {y_2 - y_1 \over x_2 - x_1 }\)

OpenStudy (lgbasallote):

yup you're right but you can cancel the negatives

Parth (parthkohli):

Yes, -3/-5 can be simplified. Cancel the negatives ;0

OpenStudy (anonymous):

thanks guys! your amazing!